OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 2(n - 3) + 5 = 3(n - 1)

OpenStudy (eyust707):

so first get rid of the parenthesis by bringing the 2 and 3 in

OpenStudy (eyust707):

once you do what do you get?

OpenStudy (anonymous):

I am confused .

OpenStudy (anonymous):

\[2(n-3)+5=3(n-1)\] 2n-6+5=3n-3 2n-1=3n-3 n=2

OpenStudy (eyust707):

ok so do you know how to distribute the number outside the ( ) into the inside?

OpenStudy (anonymous):

not really

OpenStudy (eyust707):

ok lets supposed we have 5(x+2) there is a property in math that says you are allowed to take the 5 and multiply it first by the x and then by the 2 sooo... 5(x+2) = 5x+10

OpenStudy (eyust707):

do you see how i went from 5(x+2) and made it 5x+10?

OpenStudy (anonymous):

Yes, because didn't you multiply 2 and 5?

OpenStudy (eyust707):

yea i took the 5 and multiplied it by each side of the + sign

OpenStudy (eyust707):

ok lets try one... what can we say 3(x+2) is equal to?

OpenStudy (anonymous):

3(x+2)=3x+6

OpenStudy (eyust707):

ding ding ding!!! =P

OpenStudy (eyust707):

ok so lets go back to your first equation and use that rule to solve it

OpenStudy (eyust707):

2(n - 3) + 5

OpenStudy (eyust707):

whats that

OpenStudy (anonymous):

2(n-3)+5=6n+5 ? This one is confusing me

OpenStudy (eyust707):

close!

OpenStudy (eyust707):

2(n-3)+5 = 2n-6 +5

OpenStudy (eyust707):

if you notice you have to be careful where that ( ) ends since it ends before the +5 the 2 isn't allowed to mess with the +5

OpenStudy (eyust707):

it can only be multiplied to the n-3

OpenStudy (eyust707):

do you see how I got that?

OpenStudy (anonymous):

Yesssss.

OpenStudy (eyust707):

=)

OpenStudy (eyust707):

ok what about the other side 3(n - 1) whats that equal to?

OpenStudy (anonymous):

3n-1

OpenStudy (eyust707):

close but if you notice the 3 touches the parentheses and theres two things inside there

OpenStudy (eyust707):

since the 3 touches the parenthesis it has to effect both things inside

OpenStudy (anonymous):

\[n=2\] \[2n-1=3n-3=3\]